.Determine the length of curve π‘₯ =

2

3

(𝑦 βˆ’ 1)

3

2 when 𝑦 ∈ [1,4]


1
Expert's answer
2021-05-11T13:06:35-0400

Given

"x=\frac{2}{3}(y-1)^\frac{3}{2}"

Length of curve is given by "L=\int_{0}^4 \sqrt{1+(\frac{dx}{dy})^2}dy"

Let us evaluate the above definite integral.

By differentiating with respect to y,


"\frac{dx}{dy} =(y-1)^{1/2}"

So, the integrand can be simplified as


"\sqrt{1+(\frac{dx}{dy})^2}=\sqrt{1+[(y-1)^{1/2}]^2}=\sqrt{y}=y^{1/2}"

We have,


"L=\int_{0}^4y^{1/2}dy=[\frac{2}{3}y^{3/2}]_0^4 \newline=\frac{2}{3}(4)^{3/2}-2/3(0)^{3/2}\newline=16/3"

Hence the length of the curve is "\frac{16}{3}" .